OpenStudy (anonymous):

absolute value 21-3x=2

OpenStudy (anonymous):

a hard one(show me step to step solution) evaluate the definite integral by making a u substitution and integrating from u(a) to u(b) integral 0 to pi/2, cosx/(4+4sinx)^3dx

Parth (parthkohli):

There are two possibilities: either 21 - 3x = 2 or 21 - 3x = -2

OpenStudy (anonymous):

hurry up been waiting 30min

OpenStudy (anonymous):

help

Parth (parthkohli):

Hi, can you solve it now?

OpenStudy (anonymous):

no its pretty hard for me

Parth (parthkohli):

If |21-3x| = 2 then 21 - 3x = 2 OR 21 - 3x = -2 Solve both equations individually.

OpenStudy (anonymous):

ohhhh